# Current-induced molecular dissociation: Topological insulators as robust reaction platforms

### v1

@@ III.3 Robustness against vacancy disorder
Figure 4: (a) and (b) Averaged matrix element of the occupation kernel $\bm{\mathcal{K}}^{\rho}_{\mathrm{L}}]_{\mathrm{a},\mathrm{a}}$, indicating antibonding level’s occupation from the left contact, for disordered graphene (a) and Kane-Mele (b) models. (c) Normalized energy integrals of the previous matrix elements (in the shown energy range) and those related with the bonding level $\bm{\mathcal{K}}^{\rho}_{\mathrm{L}}]_{\mathrm{b},\mathrm{b}}$. Here, blue and red lines indicate the graphene and Kane-Mele cases, respectively. We use open and filled points to distinguish between bonding and antibonding occupancies. (d) Normalized non-equilibrium intramolecular force, calculated from Eq. (16), for the shown data of panel c. The parameters used coincide with those of Fig. 2, with a central region of 31 unit cells (along the transport direction) each having a width of $W=\sqrt{3}(N-1)a/2$, where $N=3\times 25+2$.
An exciting property of catalytic topological substrates lies in the robustness of the surface (or edge) states against different types of disorder including, e.g., those associated with catalyst poisoning. To explore this feature, we study the effect of substrate degradation on catalytic performance, analyzing how the presence of atomic vacancies in both trivial (graphene) and topological (Kane-Mele model) substrates impacts molecular stability. In order to do so, we incorporate random carbon vacancies along the central system, while keeping pristine contacts. In our simulations, to stabilize the excess charge introduced by the vacancies, we passivate the nanoribbon by removing any carbon atoms that remain connected to the lattice through a single (nearest neighbor) hopping element.
@@
In Figs. 4(a) and (b) we show the antibonding occupation kernel $\bm{\mathcal{K}}^{\rho}_{\mathrm{L}}]_{\mathrm{a},\mathrm{a}}$ related to the left contact, [^2] Although not shown, the averaged bonding occupation kernel $\bm{\mathcal{K}}^{\rho}_{\mathrm{L}}]_{\mathrm{b},\mathrm{b}}$ develops a similar behavior as compared with the antibonding kernel. within a reduced energy window around the antibonding level, and averaged over 1000 independent disorder configurations. The computation of the average occupancies is motivated by the physical scenario in which a large number of molecules are deposited around one of the edges of the substrate, far from the contacts, in such a way that interactions between them can be neglected (i.e., low molecular concentration). Consequently, each vacancy configuration describes the substrate environment around a single molecule. In the case of graphene (panel a), we observe a significant reduction in the intensity of the occupation kernel, even for low vacancy concentrations such as 0.54 %. This reduction is accompanied by a broadening of the peak, mainly due to a random shift of the antibonding energy level around its original value for the pristine sample. These effects can be attributed to a high sensibility of the substrate wavefunction to the particular distribution of the vacancies. Even at a fixed concentration of vacancies, the energy position, height and width of the kernel peak strongly fluctuates between different samples, yielding high standard deviation values related with their averages. In contrast, the Kane-Mele model in Fig. 4(b) shows a more robust behavior: although there is a decrease in the intensity with increasing vacancy concentration, this reduction is much less pronounced compared to graphene samples. Moreover, the position and width of the average peaks remain nearly unchanged, indicating the robustness of the topological edge states to disorder.
Figure 4(c) summarizes the above analysis, showing the energy-integrated occupation kernels for both bonding and antibonding levels, normalized by the corresponding values of the pristine sample. Although these integrals were taken in a reduced energy range around the kernel peaks, they still capture the main contribution to the occupancies. Therefore, their normalized values reflect the sensitivity of the molecular occupation to disorder. For the Kane-Mele substrate, we observe an almost linear reduction, reaching 80% and 85% of the pristine bonding and antibonding values at 2.26% vacancy concentration, respectively. In contrast, for graphene, the decrease is more abrupt, showing a sublinear behavior to about 63% and 46% of the pristine bonding and antibonding values, respectively, at the same vacancy concentration.
@@ A.1 Explicit calculations
A.2 Finite $\eta$ broadening effects
@@ A.3 Electronic force
$\mathcal{K}^{F}_{\alpha,\sigma}=\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\left(\mathcal{K}^{\rho}_{\alpha,\sigma,\mathrm{b}}-\mathcal{K}^{\rho}_{\alpha,\sigma,\mathrm{a}}\right),$ (34)
@@
$F=\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}(\rho_{\mathrm{b},\sigma}-\rho_{\mathrm{a},\sigma}).$ (35)
We now analyze the non-equilibrium force in a situation where the two molecular levels fall within the bias window and compare it with the equilibrium situation where only the bonding level falls below $\mu_{0}$. When $V=0$ we obtain
$F=F^{\mathrm{eq}}=\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}\bar{\rho}_{\mathrm{b},\sigma}\,$, (36)
@@
$F=F^{\mathrm{eq}}+F^{\mathrm{ne}}=\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}\frac{\bar{\rho}_{\mathrm{b},\sigma}-\bar{\rho}_{\mathrm{a},\sigma}}{2},$ (38)
@@
$F^{\mathrm{ne}}=-\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}\frac{\bar{\rho}_{\mathrm{b},\sigma}+\bar{\rho}_{\mathrm{a},\sigma}}{2}.$ (39)
Once more, since $\mathrm{d}\gamma_{\mathrm{d}}/\mathrm{d}x$ is negative, the equilibrium force always acts to contract the molecule, while the non-equilibrium component promotes its destabilization. Additionally, for the situation observed in this work, where $\bar{\rho}_{\mathrm{a}}>\bar{\rho}_{\mathrm{b}}$, with $\bar{\rho}_{m}=\sum_{\sigma}\bar{\rho}_{m,\sigma}/2$, we obtain that the magnitude of the non-equilibrium force exceeds that of the equilibrium term.
Appendix B Molecular occupancies and substrate size

### v2

@@ III.3 Robustness against vacancy disorder
Figure 4: (a)-(d) Averaged matrix element of the occupation kernel $\bm{\mathcal{K}}^{\rho}_{\mathrm{L}}]_{m,m}$, indicating bonding (a,b), and antibonding (c,d) level’s occupation from the left contact, for disordered graphene (a,c) and Kane-Mele (b,d) models. (e) Normalized energy integrals of the previous matrix elements. Here, blue and red lines indicate graphene and Kane-Mele cases, respectively. We use open and filled points to distinguish between bonding and antibonding occupancies. (f) Normalized non-equilibrium intramolecular force, calculated from Eq. (16), for the shown data of panel e. The parameters used coincide with those of Fig. 2, with a central region of 31 unit cells (along the transport direction) each having a width of $W=\sqrt{3}(N-1)a/2$, where $N=3\times 25+2$.
An exciting property of catalytic topological substrates lies in the robustness of the surface (or edge) states against different types of disorder including, e.g., those associated with catalyst poisoning. To explore this feature, we study the effect of substrate degradation on catalytic performance, analyzing how the presence of atomic vacancies in both trivial (graphene) and topological (Kane-Mele model) substrates impacts molecular stability. In order to do so, we incorporate random carbon vacancies along the central system, while keeping pristine contacts [see Fig. 1(a)]. In our simulations, to stabilize the excess charge introduced by the vacancies, we passivate the nanoribbon by removing any carbon atoms that remain connected to the lattice through a single (nearest neighbor) hopping element.
When introducing carbon vacancies, the standard procedure [46, 47] is to remove both the onsite and all hopping terms connected to the corresponding site (including nearest- and next-nearest-neighbor hoppings). In doing so, the phase from the next-nearest-neighbor hoppings is symmetrically removed from both spin subspaces, thus preserving time-reversal symmetry.
@@
In Figs. 4(a)-(d) we show the bonding and antibonding occupation kernels $\bm{\mathcal{K}}^{\rho}_{\mathrm{L}}]_{m,m}$ related to the left contact, within a reduced energy window around the molecular level, and averaged over 1000 independent disorder configurations. The computation of the average occupancies is motivated by the physical scenario in which a large number of molecules are deposited around one of the edges of the substrate, far from the contacts, in such a way that interactions between them can be neglected (i.e., low molecular concentration). Consequently, each vacancy configuration describes the substrate environment around a single molecule. In the case of graphene (panels a and c), we observe a significant reduction in the intensity of the occupation kernels, even for low vacancy concentrations such as 0.54 %. This reduction is accompanied by a broadening of the peak, mainly due to a random shift of the molecular level around its original value for the pristine sample. These effects can be attributed to a high sensibility of the substrate wavefunction to the particular distribution of the vacancies. Even at a fixed concentration of vacancies, the energy position, height and width of the kernel peak strongly fluctuates between different samples, yielding high standard deviation values related with their averages. In contrast, the Kane-Mele model in Figs. 4(b) and (d) shows a more robust behavior: although there is a decrease in the intensity with increasing vacancy concentration, this reduction is much less pronounced compared to graphene samples. Moreover, the position and width of the average peaks remain nearly unchanged, indicating the robustness of the topological edge states to disorder.
Figure 4(e) summarizes the above analysis, showing the energy-integrated occupation kernels for both bonding and antibonding levels, normalized by the corresponding values of the pristine sample. Although these integrals were taken in a reduced energy range around the kernel peaks, they still capture the main contribution to the occupancies. Therefore, their normalized values reflect the sensitivity of the molecular occupation to disorder. For the Kane-Mele substrate, we observe an almost linear reduction, reaching 80% and 85% of the pristine bonding and antibonding values at 2.26% vacancy concentration, respectively. In contrast, for graphene, the decrease is more abrupt, showing a sublinear behavior to about 63% and 46% of the pristine bonding and antibonding values, respectively, at the same vacancy concentration.
@@ IV Conclusions
A detailed study of an ensemble of molecular configurations lies beyond the scope of the present work. Nevertheless, as we discuss in Appendix B, for the topological substrate the average behavior would approximately correspond to the individual molecular response. In contrast, for a non-topological substrate, this behavior may not be trivially predictable, since even at low concentrations the ensemble could act as an additional source of disorder.
@@ A.1 Explicit calculations
A.2 Finite $\eta$-broadening effects
@@ A.3 Electronic force
$\mathcal{K}^{F}_{\alpha,\sigma}=\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\left(\mathcal{K}^{\rho}_{\alpha,\sigma,+}-\mathcal{K}^{\rho}_{\alpha,\sigma,-}\right),$ (36)
@@
$F=\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}(\rho_{+,\sigma}-\rho_{-,\sigma})=2\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}(\rho_{+}-\rho_{-}),$ (37)
where we define $\rho_{m}=\sum_{\sigma}\rho_{m,\sigma}/2$, with $m=\{+,-\}$. We now analyze the non-equilibrium force in a situation where the two molecular levels fall within the bias window and compare it with the equilibrium situation where only the symmetric level falls below $\mu_{0}$. For $V=0$ we can approximate the equilibrium force as:
$F=F^{\mathrm{eq}}\approx\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}\bar{\rho}_{+,\sigma}\,$, (38)
@@
$F=F^{\mathrm{eq}}+F^{\mathrm{ne}}\approx\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}\frac{\bar{\rho}_{+,\sigma}-\bar{\rho}_{-,\sigma}}{2},$ (40)
@@
$F^{\mathrm{ne}}\approx-\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}x}\sum_{\sigma}\frac{\bar{\rho}_{+,\sigma}+\bar{\rho}_{-,\sigma}}{2}.$ (41)
Once more, since $\mathrm{d}\gamma_{\mathrm{d}}/\mathrm{d}x$ is negative, the equilibrium force always acts to compress the molecule, while the non-equilibrium component promotes its destabilization.
Appendix B $\theta$-dependence of molecular occupation and force
Figure 8: (a) Bonding (solid lines) and antibonding (dotted lines) occupancies for graphene (blue) and Kane-Mele (red) substrates as functions of the molecular angle $\theta$ [see Fig. 1(a)]. (b) Corresponding non-equilibrium intramolecular forces. We fix the bias voltage at $V=0.1\,\gamma_{0}$ and keep the other parameters the same as in Fig. 2.
In Fig. 8, we show the molecular occupancies (panel a) and the intramolecular force $F_{\mathrm{d}}$ (panel b) for a fixed bias as functions of the molecule’s angle $\theta$ subtended between the molecular bond direction and the transport axis, according to the inset of Fig. 1(a) and Eq. (4). In particular, the case $\theta=\pi/2$ corresponds to that of Figs. 2(b) and (c) for $V=0.1\,\gamma_{0}$.
As expected (since we consider a homonuclear dimer), both the molecular occupancies and the intramolecular force are symmetric functions of $\theta$ around $\pi/2$. These quantities exhibit minimal variation when the molecule is placed on a Kane-Mele substrate. Conversely, on a graphene substrate, a larger amplitude is observed for the bonding occupancy, resulting in a more pronounced related force. Nevertheless, this force remains positive (repulsive) across the full angular range, and is smaller than that of the Kane-Mele substrate.
The above discussion allows to draw an initial hypothesis on how the obtained results would change when considering an ensemble of molecules with different orientations and positions near the ribbon’s edge. In the limit of low molecular concentrations (negligible intermolecular interactions), and considering the low sensitivity to disorder on the dissociative force observed for topological substrates (see Section III.3), it can be anticipated that the ensemble behavior would approximately correspond to the average of the individual molecular responses on such substrates. In contrast, for a non-topological substrate, the strong dependence of the current-induced dissociative forces on disorder suggests that the overall ensemble behavior may not be trivially predictable, since even at low concentrations, the ensemble of molecules could act as an additional source of disorder.
Appendix C Molecular occupancies and substrate size